OpenStudy (anonymous):

Find h' (2) , given h(x) = f(g(x)) , f(u)= u^2 - 1, g(2)=3, and g'(2) -1

OpenStudy (anonymous):

I believe h'(x) = f' (g(x)) g'(x)

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

but g(X) is not given...and what is f(u)= u^2 - 1 for..

OpenStudy (anonymous):

\[f(u)=u^2-1\] so \[f(g(x))=g^2(x)-1\]

OpenStudy (anonymous):

how is g(x) = u..?

OpenStudy (anonymous):

\[f(\zeta)=\zeta^2-1\]

OpenStudy (anonymous):

srry..i dont get that... what is that equation...

OpenStudy (anonymous):



OpenStudy (anonymous):

what does this mean?

OpenStudy (anonymous):

it says solve

OpenStudy (anonymous):

solve the solution x=_____________

OpenStudy (anonymous):

how do u find f' g(x)?